a ball is thrown  verticaly up with a velocity u from top of a tower .if it strikes the ground with a velocity of 3u the time taken by the ball to reach the ground is
 

Grade:12th pass

1 Answers

Dhiraj Jain
6 years ago
t=g/2u As we know that s= v-u/t therefore we can get value of s which is equal to height of building by using formula v^2 -u^2 = 2as acceleration is g now by putting all values we will get result t is equal to g/2u . Hope it would be helpful for others.